How do you put (7,-2) m=8/7 into slope intercept form?
Answer by stanbon(75887)

How do you put (7,-2) m=8/7 into slope intercept form?
---------
Form: y = mx + b
Solve for "b":
-2 = (8/7)7 + b
-2 = 8 + b
b = -10
------
Equation:
y = (8/7)x - 10
-------------------
Cheers,
Stan H.
===================